# Advancing Super-Resolution in Neural Radiance Fields via Variational Diffusion Strategies

### v1

@@
Advancing Super-Resolution in Neural Radiance Fields via Variational Diffusion Strategies
Shrey Vishen Monta Vista High School
Cupertino, California
shrey.vishen@gmail.com
   Jatin Sarabu1, Chinmay Bharathulwar1 Bellarmine College Preparatory
Santa Clara, California
jatinsarabu13@gmail.com
John P Stevens High School
Edison, New Jersey
cbharathulwar3@gmail.com
   Rithwick Lakshmanan2, Vishnu Srinivas2 Pleasant Valley High School
Bettendorf, Iowa
rithwick.laks@gmail.com
Foothill High School
Pleasanton, California
vishnuvishaks@gmail.com
(July 2024)
I Abstract
@@
We present a novel method for diffusion-guided frameworks for view-consistent super-resolution (SR) in neural rendering. Our approach leverages existing 2D SR models in conjunction with advanced techniques such as Variational Score Distilling (VSD) and a LoRA fine-tuning helper, with spatial training to significantly boost the quality and consistency of upscaled 2D images compared to the previous methods in the literature, such as Renoised Score Distillation (RSD) proposed in DiSR-NeRF (1), or SDS proposed in DreamFusion. The VSD score facilitates precise fine-tuning of SR models, resulting in high-quality, view-consistent images. To address the common challenge of inconsistencies among independent SR 2D images, we integrate Iterative 3D Synchronization (I3DS) from the DiSR-NeRF framework (11). Our quantitative benchmarks and qualitative results on the LLFF dataset demonstrate the superior performance of our system compared to existing methods such as DiSR-NeRF. All our code is available at https://github.com/shreyvish5678/SR-NeRF-with-Variational-Diffusion-Strategies
Relevant Keywords: Diffusion Models, Super Resolution, Neural Radiance Fields (NeRF), Variational Score Distilling (VSD), Renoised Score Distillation (RSD), Score Distillation Sampling (SDS)
II Introduction
@@
Neural Radiance Fields (NeRFs) have revolutionized 3D scene rendering from 2D images, significantly impacting applications such as 3D reconstruction and virtual reality (2). NeRFs use continuous volumetric functions optimized by neural networks to synthesize high-fidelity views (28)(29)(30). However, scaling NeRF for super-resolution, maintaining view consistency, and managing high-dimensional data remains challenging (18)(19)(20)(31). Somewhat recent advancements such as Mip-NeRF 360 (24) and TensoRF have addressed some of these challenges with NeRF, especially with unbounded scene rendering and tensor decomposition. Our work focuses specifically on enhancing super-resolution capabilities within the NeRF framework.
@@
Previous enhancements to NeRF, like Score Distillation Sampling (SDS), have struggled with issues such as over-smoothing and computational inefficiency, limiting their ability to capture fine details. To address these problems, we introduce Variational Score Distillation (VSD), which models 3D scene parameters as probabilistic distributions rather than fixed values. This approach improves scene representation by leveraging diffusion models (5)(6) and incorporates low-rank adaptation (LoRA) for efficient fine-tuning of pre-trained models.
@@
Our extensive experiments show that VSD significantly outperforms SDS and RSD in generating detailed and photorealistic NeRFs, enhancing visual quality and computational efficiency. Results from datasets like LLFF confirm these improvements. Additionally, we provide a detailed ablation study on our system’s components, including LoRA-based fine-tuning (7) and hierarchical sampling strategies(13).
@@
Our approach advances NeRF-based rendering, offering a new standard for high-resolution 3D scene generation with broad applications in entertainment, gaming, scientific visualization, and architectural design (15)(16)(17).
III Methodolgy
III-A Pre-Requisites
III-A1 Latent Encoding and Residual Learning
@@
Latent encoding starts by extracting 2D projections from a lower-resolution Neural Radiance Field (NeRF). These 2D views are then passed through an encoder, transforming them into latent space representations that capture key features in a compressed form (3).
@@
$x_{0}^{\prime}=x_{0}+h_{\theta}$
III-A2 Forward Diffusion Process
Once the latent vectors, including the residual latents, are generated, they undergo a forward diffusion process (4). Noise is gradually added to the latent vectors over timesteps, governed by the equation:
$x_{t}=\sqrt{\alpha_{t}}x_{0}+\sqrt{1-\alpha_{t}}\epsilon,\quad\epsilon\sim\mathcal{N}(0,I)$
@@ III-A2 Forward Diffusion Process
In the subsequent denoising step, the noisy latents are refined to produce high-quality, view-consistent images (5)(6). The overall goal of this process is to prepare the latent vectors for reconstruction, reversing the noise added during forward diffusion.
III-A3 UNet-Based Prediction
In our model, the prediction of the final high-quality image relies heavily on the use of UNet architectures. The UNet models are employed to process the noisy latent vectors generated during the forward diffusion process and to reconstruct these into cleaner, more accurate representations of the original scene.
Pre-trained UNet Model and LoRA-adapted Model
@@ III-A3 UNet-Based Prediction
$\epsilon_{\phi}=f_{\phi}(\mathbf{x}_{t},\mathbf{t},\mathbf{y}^{\mathbf{c}},\mathbf{I}_{LR}),$
@@
The noisy latent vector $\mathbf{x}_{t}$ is also processed by a Fine-Tuned UNet model, which includes learnable Low-rank Adaptation (LoRA) parameters (7). These allow the model to efficiently adapt for high-quality image reconstruction. Unlike the frozen pre-trained model, the Fine-Tuned UNet is trained during this stage, with LoRA parameters enabling targeted adjustments to the network layers.
The Fine-Tuned UNet takes the same inputs: $\mathbf{x}_{t}$, time embeddings $\mathbf{t}$, text embeddings $\mathbf{y}$, and low-resolution image $\mathbf{I}_{LR}$, plus class labels $\mathbf{c}$ to focus on task-specific features (8). Its prediction is:
$\epsilon_{\varphi}=f_{\varphi}(\mathbf{x}_{t},\mathbf{t},\mathbf{y},\mathbf{I}_{LR},\mathbf{c}),$
@@
Comparison and Loss Calculation
To quantify the improvement in image quality provided by the Fine-Tuned UNet, we compute the Variational Score Distillation (VSD) loss, which measures the difference between the predictions of the pre-trained and fine-tuned models. We used an L1 variation of the VSD loss proposed in Prolific Dreamer (12), defined as:
$\mathcal{L}_{\text{VSD}}(\theta)$ $=\mathbb{E}_{t,\epsilon,\mathbf{c}}\Bigg{[}\omega(t)\cdot\Bigg{\|}\epsilon_{\phi}-\epsilon_{\varphi}\Bigg{\|}\Bigg{]}$
@@
This loss is then backpropagated through the network, specifically targeting the LoRA parameters in the Fine-Tuned UNet. By minimizing this loss, the model learns to generate high-quality latent representations that closely match the ideal outputs while incorporating task-specific adjustments through the class labels. Then we obtrain our new residual latents:
$\theta_{\text{new}}=\theta_{\text{old}}-\eta\nabla_{\theta}\mathcal{L}_{\text{VSD}}(\theta),$
@@
$\mathcal{L}_{\text{Diff}}(\theta)$ $=\mathbb{E}_{t,\epsilon,\mathbf{c}}\Bigg{[}f_{\varphi}(\mathbf{x}_{t},t,\mathbf{c},\mathbf{I}_{LR},\mathbf{y})-\epsilon\Bigg{]}^{2}$
@@
$\varphi_{\text{new}}=\varphi_{\text{old}}-\eta\nabla_{\varphi}\mathcal{L}_{\text{Diff}}(\theta),$
@@
This dual UNet approach, leveraging both pre-trained and fine-tuned models, ensures that the latent prediction process is both accurate and adaptable, ultimately contributing to the generation of high-resolution, view-consistent images that surpass the quality of those produced by existing methods (8)(9).
III-A4 Low-rank Adaptation (LoRA) Fine-Tuning
Low-rank Adaptation (LoRA) is a powerful technique designed to fine-tune pre-trained neural networks efficiently, particularly in scenarios where extensive retraining or structural modifications are impractical (7). In the context of Neural Radiance Fields (NeRFs), LoRA introduces trainable low-rank matrices into selected layers of the network, allowing for effective adaptation to new datasets or specific rendering tasks with minimal computational overhead.
Concept and Integration
@@ III-A4 Low-rank Adaptation (LoRA) Fine-Tuning
$\frac{\partial\mathcal{L}}{\partial A}=\frac{\partial\mathcal{L}}{\partial W^{\prime}}B^{T},\quad\frac{\partial\mathcal{L}}{\partial B}=A^{T}\frac{\partial\mathcal{L}}{\partial W^{\prime}}$
@@
III-B Lower Resolution NeRF Generation
The first step in our pipeline involves constructing a low-resolution (LR) Neural Radiance Field (NeRF) from an LR image dataset. Using Nvidia’s InstantNGP framework (26), we efficiently build this NeRF model. Afterward, the LR images are upsampled to match the resolution of the target super-resolution (SR) images, establishing a baseline for our iterative training process (14)(27).
III-C SR Training Process
The training process builds upon recent NeRF advancements, utilizing insights from a variety of volumetric rendering techniques (21,22) and neural field applications (23). The training loop operates iteratively until convergence, following a multi-step approach:
1. Random Render Sampling: A random LR image is rendered from the NeRF and used as input for encoding.This image is then taken and interpolated to 4x it’s size, and then converted into a latent. Both the image and latent are passed.
@@ III-C SR Training Process
This process repeats until the model converges, producing high-resolution, view-consistent NeRF outputs that outperform existing methods (10)(32), as we can see in Algorithm 1
@@
III-D Iterative 3D Sync Process
Our Iterative 3D Synchronization (I3DS) approach addresses limitations observed in initial experiments, where applying SDS directly on NeRF renders resulted in blurred details and convergence issues. We propose I3DS to improve this by decoupling the upscaling and NeRF synchronization processes into two alternating stages.
@@ IV-A Statistical Analysis
IV-B Comparison with Existing Methods
@@
TABLE I: Comparison of Proposed Method with Existing Methods
@@
No Changes (Plain RSD) 0.14955 4.9831 3.9827
With SDS 0.15871 5.6671 3.5288
With VSD + LoRA Spaced 0.15233 4.4573 4.0261
With VSD + LoRA 0.15498 4.6122 3.9984
@@ V Conclusion
VI Acknowledgments
We would like to thank Ayaan Haque and Michael Lutz for their guidance throughout the research project.

### v2

@@
Abstract
@@
We present a novel method for diffusion-guided frameworks for view-consistent super-resolution (SR) in neural rendering. Our approach leverages existing 2D SR models in conjunction with advanced techniques such as Variational Score Distilling (VSD) and a LoRA fine-tuning helper, with spatial training to significantly boost the quality and consistency of upscaled 2D images compared to the previous methods in the literature, such as Renoised Score Distillation (RSD) proposed in DiSR-NeRF [7], or SDS proposed in DreamFusion. The VSD score facilitates precise fine-tuning of SR models, resulting in high-quality, view-consistent images. To address the common challenge of inconsistencies among independent SR 2D images, we integrate Iterative 3D Synchronization (I3DS) from the DiSR-NeRF framework [5]. Quantitative benchmarks and qualitative results on the LLFF dataset demonstrate the superior performance of our system compared to existing methods such as DiSR-NeRF. All our code is available at https://github.com/shreyvish5678/SR-NeRF-with-Variational-Diffusion-Strategies
1 Introduction
@@
Neural Radiance Fields (NeRFs) have revolutionized 3D scene rendering from 2D images, significantly impacting applications such as 3D reconstruction and virtual reality [10]. NeRFs use continuous volumetric functions optimized by neural networks to synthesize high-fidelity views [19][30][21]. However, scaling NeRF for super-resolution, maintaining view consistency, and managing high-dimensional data remains challenging [31][32][20][27]. Somewhat recent advancements such as Mip-NeRF 360 [2] and TensoRF have addressed some of these challenges with NeRF, especially with unbounded scene rendering and tensor decomposition. Our work focuses specifically on enhancing super-resolution capabilities within the NeRF framework.
@@
Previous enhancements to NeRF, like Score Distillation Sampling (SDS), have struggled with issues such as over-smoothing and computational inefficiency, limiting their ability to capture fine details. To address these problems, we introduce Variational Score Distillation (VSD), which models 3D scene parameters as probabilistic distributions rather than fixed values. This approach improves scene representation by leveraging diffusion models [4][13] and incorporates low-rank adaptation (LoRA) for efficient fine-tuning of pre-trained models.
@@
Our extensive experiments show that VSD significantly outperforms SDS and RSD in generating detailed and photorealistic NeRFs, enhancing visual quality and computational efficiency. Results from datasets like LLFF confirm these improvements. Additionally, we provide a detailed ablation study on our system’s components, including LoRA-based fine-tuning [6] and hierarchical sampling strategies[22].
@@
Our approach advances NeRF-based rendering, offering a new standard for high-resolution 3D scene generation with broad applications in entertainment, gaming, scientific visualization, and architectural design [1][9][15].
To summarize, the contributions of this paper are as follows,
  • Using Pre-trained Stable Diffusion weights for 3D NeRF Render Upscaling
  • Utilizing Low-rank adaptation with Mixed Precision training for steering outputs
  • Using a version of VSD loss and I3DS for training and upscaling the NeRF outputs
2 Methodolgy
2.1 Pre-Requisites
2.1.1 Latent Encoding and Residual Learning
@@
Latent encoding starts by extracting 2D projections from a lower-resolution Neural Radiance Field (NeRF). These 2D views are then passed through an encoder, transforming them into latent space representations that capture key features in a compressed form [17].
@@
$x_{0}^{\prime}=x_{0}+h_{\theta}$ (1)
2.1.2 Forward Diffusion Process
Once the latent vectors, including the residual latents, are generated, they undergo a forward diffusion process [18]. Noise is gradually added to the latent vectors over timesteps, governed by the equation:
$x_{t}=\sqrt{\alpha_{t}}x_{0}+\sqrt{1-\alpha_{t}}\epsilon,\quad\epsilon\sim\mathcal{N}(0,I)$ (2)
@@ 2.1.2 Forward Diffusion Process
In the subsequent denoising step, the noisy latents are refined to produce high-quality, view-consistent images [4][13]. The overall goal of this process is to prepare the latent vectors for reconstruction, reversing the noise added during forward diffusion.
2.1.3 UNet-Based Prediction
In the model, the prediction of the final high-quality image relies heavily on the use of UNet architectures. The UNet models are employed to process the noisy latent vectors generated during the forward diffusion process and to reconstruct these into cleaner, more accurate representations of the original scene.
@@ 2.1.3 UNet-Based Prediction
$\epsilon_{\phi}=f_{\phi}(\mathbf{x}_{t},\mathbf{t},\mathbf{y}^{\mathbf{c}},\mathbf{I}_{LR})$ (3)
@@
The noisy latent vector $\mathbf{x}_{t}$ is also processed by a Fine-Tuned UNet model, which includes learnable Low-rank Adaptation (LoRA) parameters [6]. These allow the model to efficiently adapt for high-quality image reconstruction. Unlike the frozen pre-trained model, the Fine-Tuned UNet is trained during this stage, with LoRA parameters enabling targeted adjustments to the network layers.
The Fine-Tuned UNet takes the same inputs: $\mathbf{x}_{t}$, time embeddings $\mathbf{t}$, text embeddings $\mathbf{y}$, and low-resolution image $\mathbf{I}_{LR}$, plus class labels $\mathbf{c}$ to focus on task-specific features [8]. Its prediction is:
$\epsilon_{\varphi}=f_{\varphi}(\mathbf{x}_{t},\mathbf{t},\mathbf{y},\mathbf{I}_{LR},\mathbf{c})$ (4)
@@
To quantify the improvement in image quality provided by the Fine-Tuned UNet, we compute the Variational Score Distillation (VSD) loss, which measures the difference between the predictions of the pre-trained and fine-tuned models. We used an L1 variation of the VSD loss proposed in Prolific Dreamer [26], defined as:
$\mathcal{L}_{\text{VSD}}(\theta)=\mathbb{E}_{t,\epsilon,\mathbf{c}}\left[\omega(t)\cdot\left\|\epsilon_{\phi}-\epsilon_{\varphi}\right\|\right]$ (5)
@@
This loss is then backpropagated through the network, specifically targeting the LoRA parameters in the Fine-Tuned UNet. By minimizing this loss, the model learns to generate high-quality latent representations that closely match the ideal outputs while incorporating task-specific adjustments through the class labels. Then we obtrain the new residual latents:
$\theta_{\text{new}}=\theta_{\text{old}}-\eta\nabla_{\theta}\mathcal{L}_{\text{VSD}}(\theta)$ (6)
@@
$\mathcal{L}_{\text{Diff}}(\theta)=\mathbb{E}_{t,\epsilon,\mathbf{c}}\left[\left(f_{\varphi}(\mathbf{x}_{t},t,\mathbf{c},\mathbf{I}_{LR},\mathbf{y})-\epsilon\right)^{2}\right]$ (7)
@@
$\varphi_{\text{new}}=\varphi_{\text{old}}-\eta\nabla_{\varphi}\mathcal{L}_{\text{Diff}}(\theta)$ (8)
@@
This dual UNet approach, leveraging both pre-trained and fine-tuned models, ensures that the latent prediction process is both accurate and adaptable, ultimately contributing to the generation of high-resolution, view-consistent images that surpass the quality of those produced by existing methods [8][25].
2.2 Low-rank Adaptation (LoRA) Fine-Tuning
Low-rank Adaptation (LoRA) is a powerful technique designed to fine-tune pre-trained neural networks efficiently, particularly in scenarios where extensive retraining or structural modifications are impractical [6]. In the context of Neural Radiance Fields (NeRFs), LoRA introduces trainable low-rank matrices into selected layers of the network, allowing for effective adaptation to new datasets or specific rendering tasks with minimal computational overhead.
@@ 2.2 Low-rank Adaptation (LoRA) Fine-Tuning
$\frac{\partial\mathcal{L}}{\partial A}=\frac{\partial\mathcal{L}}{\partial W^{\prime}}B^{T},\quad\frac{\partial\mathcal{L}}{\partial B}=A^{T}\frac{\partial\mathcal{L}}{\partial W^{\prime}}$ (10)
@@
2.3 Lower Resolution NeRF Generation
The first step in our pipeline involves constructing a low-resolution (LR) Neural Radiance Field (NeRF) from an LR image dataset. Using Nvidia’s InstantNGP framework [12], we efficiently build this NeRF model. Afterward, the LR images are upsampled to match the resolution of the target super-resolution (SR) images, establishing a baseline for the iterative training process [14][16][3].
2.4 SR Training Process
The training process builds upon recent NeRF advancements, utilizing insights from a variety of volumetric rendering techniques [23][29] and neural field applications [28]. The training loop operates iteratively until convergence, following a multi-step approach:
1. Random Render Sampling: A random LR image is rendered from the NeRF and used as input for encoding. This image is then taken and interpolated to 4x it’s size, and then converted into a latent. Both the image and latent are passed. A text prompt is passed as well which describes the NeRF content along with from what orientation the image should look like, to condition the model for the desired output.
@@ 2.4 SR Training Process
This process repeats until the model converges, producing high-resolution, view-consistent NeRF outputs that outperform existing methods [11][24], as we can see in Algorithm 1
@@
2.5 Iterative 3D Sync Process
The Iterative 3D Synchronization (I3DS) approach addresses limitations observed in initial experiments, where applying SDS directly on NeRF renders resulted in blurred details and convergence issues. We add I3DS to the pipeline to improve this by decoupling the upscaling and NeRF synchronization processes into two alternating stages.
@@ 3.1 Statistical Analysis
3.2 Comparison with Existing Methods
@@
Table 1: Comparison of Proposed Method with Existing Methods
@@
No Changes (Plain RSD) 0.1496 4.983 3.983
With SDS 0.1587 5.667 3.529
With VSD + LoRA Spaced 0.1523 4.457 4.026
With VSD + LoRA 0.1550 4.612 3.998